OpenStudy (anonymous):

Trevor is tiling his bathroom floor, which has an area that is represented as 175r5 square inches. Each tile has an area of sqrt(64r^17) . The total number of tiles used can be represented by the expression below. 175r^5/sqrt(64r^17) Simplify the expression for the total number of tiles used. Show your work.

OpenStudy (anonymous):

lets work it one part at a time. I want to do the denominator first

OpenStudy (anonymous):

\[\sqrt{64r^{17}}=\sqrt{64}\sqrt{r^{17}}=8(r^{17})^\frac{ 1 }{ 2 }=8r^{ \frac{ 17 }{ 2 }}\]

OpenStudy (anonymous):

Does that make sense to you? Any questions on this part?

OpenStudy (anonymous):

nope. i got that part.

OpenStudy (anonymous):

Now put them together. \[\frac{ 175r^5 }{ 8r^{\frac{ 17 }{ 2 } }}=\frac{ 175 }{ 8 }\frac{ r^5 }{ r^{\frac{ 17 }{ 2 }} }\]

OpenStudy (anonymous):

\[\frac{ x^n }{ x^m }=x^{(n-m)}\]

OpenStudy (anonymous):

Any questions so far?

OpenStudy (anonymous):

using the formula I gave you. I want you to simplify the r's. let me know what you get

OpenStudy (anonymous):

how would yu simplify the 17/2?

OpenStudy (anonymous):

You can't that is the simplest fraction. but you need to cancel with r^5 in the numerator. Use the formula I gave x^(n-m), where n and m are real numbers

OpenStudy (anonymous):

oh okay. i think i got it

OpenStudy (anonymous):

If you post your answer I'll check it.